### v1

@@ 1 Introduction
The minimum 20% post-CE fraction already has one critical implication for our understanding of PNe – it is likely that some stars which our textbooks would say should produce an observable PN, in fact, do not. Simply put, based on the main sequence binary fraction and period distribution, one would not expect as many as 20% of all low- and intermediate-mass stars to experience a CE phase, let alone survive one without merging. Clearly, with such a large uncertainty on the true post-CE binary fraction in PNe (20–80%), it is similarly uncertain just how large the discrepancy between the main sequence and PN phase really is. However, there are indications that it may indeed be significant. Population synthesis studies have found that perhaps around one fifth of expected PN progenitors due not lead to observable PNe (potentially consistent with only binaries leading to PNe; Moe & De Marco, 2006), while as many as 50% of central stars exhibit properties that are inconsistent with single-star evolutionary tracks (Weidmann et al., 2020).
@@
As a result of conservation of angular momentum during the ejection, the CE is always expected to be preferentially deposited in the orbital plane of the binary thus leading to nebular symmetry axes which lay perpendicular to this plane. While only nine PNe have been the subject of sufficiently detailed study to constrain both their nebular inclinations and central binary orbital inclinations, the correlation is striking and entirely consistent with the CE shaping each and every one. Indeed, the probability of finding such an alignment in these PNe by chance has been shown to be less than one in a million (Munday et al., 2020). Nonetheless, there is significant morphological diversity, even among this sample, meaning that, while the CE has clearly played a role in their shaping, the exact processes at work may vary appreciably between (Hillwig et al., 2016).
@@
Thus, it now seems beyond doubt that CE evolution is a critical aspect of PN formation and evolution, even though the precise details of this are still the subject of active research (Jones & Boffin, 2017; Boffin & Jones, 2019). But, perhaps of greater interest for the wider astronomical community, we are now in a position to use post-CE PNe to try and understand the CE phase itself. The nebular material in these PNe is thought to comprise the ejected CE, offering a unique opportunity to study the ejecta directly (CE ejecta can also be observed in luminous red novae, but these represent “failed” CE events, where the stars merged before fully ejecting the donor’s envelope). Furthermore, the short-lived nature of PNe ($\tau\sim$30 kyr) means that post-CE central stars of PNe have not yet had time to adjust following the CE ejection. As such, post-CE PNe are an invaluable tool with which to study the CE phase leading to potentially interesting conclusions for a wide-range of other post-CE phenomena.
@@ 2 Pre-CE mass transfer
It is important to note that no inflation is observed in the MS companions of more evolved post-CE binaries (indeed these systems are excellent probes of the mass-radius relationship of late-type MS stars; Parsons et al., 2018). This might, at first, appear puzzling when one takes into account that the thermal timescales for these stars are too long for them to have relaxed back to a normal state. However, models predict that for intermediate values of accretion efficiency and rate, the response of fully convective stars is to become partially convective (Prialnik & Livio, 1985), and it is thus the convective layer formed in the outer envelope which expands rather than the star as a whole. The relaxation of this convective layer can then occur on a much shorter timescale (Jones et al., 2022).
@@ 6 Pre-PNe and the link to luminous red novae
Intriguingly, the ejecta properties of luminous red novae – a class of transients associated with CE mergers – are strikingly similar to those of pre-PNe (Kamiński et al., 2018), offering a potential explanation for the dearth of surviving post-CE binaries in pre-PNe. Although, it must be highlighted that much progress has been made linking the morphologies of pre-PNe with wide stellar companions (Decin et al., 2020), several of which are known in both pre-PNe (e.g., Sánchez Contreras et al., 2004) and PNe (e.g., Jones et al., 2017).

### v3

@@ 1 Introduction
The minimum 20% post-CE fraction already has one critical implication for our understanding of PNe – it is likely that some stars which our textbooks would say should produce an observable PN, in fact, do not. Simply put, based on the main sequence binary fraction and period distribution, one would not expect as many as 20% of all low- and intermediate-mass stars to experience a CE phase, let alone survive one without merging. The implication being that, if binaries appear over-abundant, then some single stars previously believed to be PN progenitors must not lead to observable PNe. Clearly, with such a large uncertainty on the true post-CE binary fraction in PNe (20–80%), it is similarly uncertain just what fraction of these single-star PN progenitors will not lead to PNe. However, there are indications that it may indeed be significant. Population synthesis studies have found that perhaps around one fifth of expected PN progenitors due not lead to observable PNe (potentially consistent with only binaries leading to PNe; Moe & De Marco, 2006), while as many as 50% of central stars exhibit properties that are inconsistent with single-star evolutionary tracks (Weidmann et al., 2020).
@@
As a result of conservation of angular momentum during the ejection, the CE is always expected to be preferentially deposited in the orbital plane of the binary thus leading to nebular symmetry axes which lay perpendicular to this plane. While only nine PNe have been the subject of sufficiently detailed study to constrain both their nebular inclinations and central binary orbital inclinations, the correlation is striking and entirely consistent with the CE shaping each and every one (Hillwig et al., 2016). Indeed, the probability of finding such an alignment in these PNe by chance has been shown to be less than one in a million (Munday et al., 2020). Nonetheless, there is significant morphological diversity, even among this sample, meaning that, while the CE has clearly played a role in their shaping, the exact processes at work may vary appreciably (Hillwig et al., 2016).
@@
Thus, it now seems beyond doubt that CE evolution is a critical aspect of PN formation and evolution, even though the precise details of this are still the subject of active research (Jones & Boffin, 2017; Boffin & Jones, 2019). But, perhaps of greater interest for the wider astronomical community, we are now in a position to use post-CE PNe to try and understand the CE phase itself. The nebular material in these PNe is thought to comprise the ejected CE, offering a unique opportunity to study the ejecta directly (CE ejecta can also be observed in luminous red novae, but these represent “failed” CE events, where the stars merged before fully ejecting the donor’s envelope). Furthermore, the short-lived nature of PNe ($\tau\sim$30 kyr) means that post-CE central stars of PNe have not yet had time to adjust their chemical and physical properties following the CE ejection. As such, post-CE PNe are an invaluable tool with which to study the CE phase leading to potentially interesting conclusions for a wide-range of other post-CE phenomena.
@@ 2 Pre-CE mass transfer
It is important to note that there is little to no inflation observed in the MS companions of more evolved post-CE binaries (indeed these systems are excellent probes of the mass-radius relationship of late-type MS stars; Parsons et al., 2018). This might, at first, appear puzzling when one takes into account that the thermal timescales for these stars are too long for them to have relaxed back to a normal state. However, models predict that for intermediate values of accretion efficiency and rate, the response of fully convective stars is to become partially convective (Prialnik & Livio, 1985), and it is thus the convective layer formed in the outer envelope which expands rather than the star as a whole. The relaxation of this convective layer can then occur on a much shorter timescale (Jones et al., 2022).
@@ 6 Pre-PNe and the link to luminous red novae
Intriguingly, the ejecta properties of luminous red novae – a class of transients associated with CE mergers – are strikingly similar to those of pre-PNe (Kamiński et al., 2018), offering a potential explanation for the dearth of surviving post-CE binaries in pre-PNe. However, it must be highlighted that much progress has been made linking the morphologies of pre-PNe with wide stellar companions (Decin et al., 2020), several of which are known in both pre-PNe (e.g., Sánchez Contreras et al., 2004) and PNe (e.g., Jones et al., 2017).